OpenStudy (anonymous):

When solving for the y-intercept, what do i do when the equation has fractions in it? For example: Q:8=12/11*5+b Do i get them so i can divide them by the same thing?(the fractions)

OpenStudy (anonymous):

Or should i divide both sides by 12/11?

OpenStudy (anonymous):

\[8=\frac{12}{11}*5+b\] Is this what you have?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Okay times both sides of the equation by 11 to get rid of the fraction

OpenStudy (anonymous):

What? like 11*12/11*11?

OpenStudy (king):

b=28/11

OpenStudy (anonymous):

Dude! Don't just GIVE me the answer! i wanted to learn how to do this!

OpenStudy (anonymous):

Ok new example.

OpenStudy (anonymous):

\[8*11=\frac{12*11}{11}*5+11*b\] like this, see the 11 in the denominator will cancel

OpenStudy (king):

oh!sorry!shall i post the solution and explain it to u?

OpenStudy (anonymous):

King, ill get a new example and ill try it :) Its ok.

OpenStudy (anonymous):

The denominator is the bottom one, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, so it becomes just 11*12?

OpenStudy (king):

sorry once again!

OpenStudy (anonymous):

No, thats ok :D

OpenStudy (anonymous):

yes 11*12*5 don't forget the 5

OpenStudy (anonymous):

Ok so now its 8*11*12*5+11b?

OpenStudy (king):

yup inopeki thats rite!

OpenStudy (king):

no there is no 8!8 is on the LHS

OpenStudy (anonymous):

Wait, isnt it 8*11=11*12*5+11b

OpenStudy (anonymous):

correct inopeki

OpenStudy (anonymous):

LHS?

OpenStudy (anonymous):

left hand side of the equation

OpenStudy (anonymous):

Ah

OpenStudy (king):

ure 2nd equation is rite.

OpenStudy (anonymous):

Now divide both sides by 11?

OpenStudy (anonymous):

getting 8=12*5+b?

OpenStudy (king):

yup !

OpenStudy (anonymous):

Multiply 12 by 5 getting 60 so now we have 8=60+b

OpenStudy (anonymous):

hold on you will have 8*11=12*5+11*b

OpenStudy (king):

now subtract both sides by 60

OpenStudy (anonymous):

Take both sides minus 60 and we have the answer! -52=b?

OpenStudy (anonymous):

Whaaat?

OpenStudy (king):

oh !!!!!!!!answer is b=-52!sry!

OpenStudy (anonymous):

Its ok, but your answer was right too, according to khanacademy :)

OpenStudy (anonymous):

Ill try another one

OpenStudy (anonymous):

5=-5/6*-9+b

OpenStudy (king):

u solve it!

OpenStudy (king):

and tell me ure answer i will tell if it matches with my answer!

OpenStudy (anonymous):

Multiply both sides of the fraction by -5 getting 5=-5*5/5*6*-9+b The 5s cancel out getting us to 5=-5*5/6*-9+b, simplify: 5=-25/6*-9+b. Am i right so far?

OpenStudy (king):

hmm.i dont get t i can tell u sumthin else to do but multiply by 5 on both sides of the equation

OpenStudy (anonymous):

Oh right!

OpenStudy (king):

then u will get?

OpenStudy (anonymous):

5=-5*5/5*6*-9*5+5b?

OpenStudy (king):

even the 5on the left hand side should be multiplied by 5 so it will bcome 25....

OpenStudy (anonymous):

Oh right, 5*5=-5*5/5*6*-9*5+5b?

OpenStudy (king):

yup so now cancel 5's...............

OpenStudy (anonymous):

Is the equation you're doing \[5=\frac{-5}{6}*-9+b\]?

OpenStudy (king):

yup i think so !

OpenStudy (anonymous):

The whole fraction is negative, but yeah

OpenStudy (anonymous):

why did you times both side by 5, it should be 6 if you want to get rid of the fraction

OpenStudy (anonymous):

Oh right

OpenStudy (anonymous):

5*6=-5*6/6*6*-9*6+6b?

OpenStudy (anonymous):

is the fraction times by -9 or just minus 9?

OpenStudy (king):

u shouldnt multiply -9 by 6

OpenStudy (anonymous):

Simplified: 30=-30/36*-9+6b

OpenStudy (anonymous):

so yes king is right, typo

OpenStudy (anonymous):

Yeah its -9 multiplied by the fraction, so we dont multiply it?

OpenStudy (anonymous):

yes don't times it

OpenStudy (anonymous):

sorry guys i just can't type well tonight :)

OpenStudy (king):

-9 is include in the fraction -5/6

OpenStudy (anonymous):

so inopeki what do you have for your equation?

OpenStudy (anonymous):

30=-30/36*-9+6b

OpenStudy (anonymous):

close, remember we times the fraction by 6 to get rid of it. so we have 30=-5*-9+6b

OpenStudy (anonymous):

How did you get -5?

OpenStudy (anonymous):

OOOH

OpenStudy (anonymous):

bacause the fraction was 5*6/6*6? The sixes cancel out!

OpenStudy (anonymous):

yes but you don't need to times the denominator by 6 as well. only time each term once!

OpenStudy (anonymous):

Oh right

OpenStudy (anonymous):

So it was -5*6/6

OpenStudy (anonymous):

cause the fraction is one term

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

and -5*6/6=-5

OpenStudy (anonymous):

perfect now put this into the equation and you get 30=-5*-9+6b which has no fraction in it at all :)

OpenStudy (anonymous):

30=-5*-9+6b Simplified: 30=45+6b

OpenStudy (king):

whts ure answer inopeki?

OpenStudy (anonymous):

both sides minus 45=6b

OpenStudy (anonymous):

so -15=6b

OpenStudy (anonymous):

So -15/6?

OpenStudy (anonymous):

perfect!!!!!

OpenStudy (anonymous):

YES!

OpenStudy (anonymous):

That was wrong :(

OpenStudy (anonymous):

How do i simplify it?

OpenStudy (anonymous):

\[\frac{-15}{6}=\frac{-15/3}{6/3}=\frac{-5}{2}\]